# Cluster algebras and quantum cohomology rings: A-type

### v1

@@ 4.4 Behavior of one-dimensional $S$-orbits under quiver mutations
$\varphi:\mathcal{Z}^{1}\rightarrow(\mathcal{Z}^{\prime})^{1}.$ (4.21)
@@
4.5 Behavior of weights $\mathbf{w}(\tau,\sigma)$ under quiver mutations
@@
The main goal of this subsection is to investigate the behavior of weights of flags defined in (4.3) under quiver mutations.
@@ 4.5.1 Grassmannian case
Let $\sigma_{1}=I=\{i_{1}<\cdots<i_{r}\},\sigma_{2}=J=\{j_{1}<\cdots<j_{r}\}$ be two $S$-fixed points in $Gr(r,N)$ which without loss of generality satisfy $i_{1}\neq j_{1},i_{k}=j_{k}$ for $k\neq 1$. Let $\tau$ be the 1-dimensional $S$-orbit connecting the two points whose matrix representation is
$\tau=\begin{bmatrix}\mathbf{0}&a&\mathbf{0}&b&\mathbf{0}\\ \mathbf{0}&\mathbf{0}&\mathbf{0}&\mathbf{0}&B\end{bmatrix}$ (4.22)
@@
$\mathbf{w}(\tau,\sigma_{1})=c_{1}^{S}(T_{\sigma_{1}}\tau)=\lambda_{i_{1}}-\lambda_{j_{1}},\,\mathbf{w}(\tau,\sigma_{2})=c_{1}^{S}(T_{\sigma_{2}}\tau)=\lambda_{j_{1}}-\lambda_{i_{1}}.$ (4.24)
@@
$\begin{bmatrix}\mathbf{0}&\mathbf{0}\\ b&\mathbf{0}\\ \mathbf{0}&\mathbf{0}\\ a&\mathbf{0}\\ \mathbf{0}&C\end{bmatrix}$ (4.25)
@@
$\mathbf{w}(\varphi(\tau_{1}),\varphi(\sigma))=\lambda_{i_{1}}-\lambda_{j_{1}},\,\mathbf{w}(\varphi(\tau),\varphi(\sigma_{2}))=\lambda_{j_{1}}-\lambda_{i_{1}}.$ (4.27)
Compare equations (4.24) and (4.27), and one can find
$\mathbf{w}(\tau,\sigma_{1})=\mathbf{w}(\varphi(\tau_{1}),\varphi(\sigma)),\,\,\mathbf{w}(\tau,\sigma_{2})=\mathbf{w}(\varphi(\tau),\varphi(\sigma_{2}))\,.$ (4.28)
@@ 4.5.2 General case
$\mathbf{w}(\tau,\sigma)=\mathbf{w}(\varphi(\tau),\varphi(\sigma)).$ (4.29)
@@
  1. (1)
    When $I_{v_{3}}=J_{v_{3}}$, $I_{v_{4}}=J_{v_{4}}$, one matrix representative $(A_{u})_{u\in Q_{0}\backslash\{v_{f}\}}$ can be written as $A_{u},u\neq v$ are in reduced row/column echelon forms whose columns/rows numbers of pivots are $I_{u},u\neq v$. Meanwhile, the matrix $A_{v}$ can be written as the same formula as (4.22). We get the vertex $\sigma_{1}$ if $a=1,b=0$ and $\sigma_{2}$ is $a=0,b=1$. There is an induced torus action $(\mathbb{C}^{*})^{r_{v_{1}}}$ on the matrix $A_{v}$ as (4.23). Hence, we have $\mathbf{w}(\tau,\sigma_{1})=\lambda_{i_{1}}-\lambda_{j_{1}}$, and $\mathbf{w}(\tau,\sigma_{2})=\lambda_{j_{1}}-\lambda_{i_{1}}$.
    After the quiver mutation $\mu_{v}$, we then have $I_{v}^{\prime}\backslash J_{v}^{\prime}=\{j_{1}\}$ and $J_{v}^{\prime}\backslash I_{v}^{\prime}=\{i_{1}\}$. The matrices representatives $A_{u}^{\prime},\,u\neq v$ are the same with $A_{u},\,u\neq v$, but $A_{v}^{\prime}$ is changed to the formula (4.25), with the induced torus action $(\mathbb{C}^{*})^{r_{v_{1}}}$ the same with (4.26). One can easily find that the weights of the flag at the vertex $\varphi(\sigma_{1})$ and $\varphi(\sigma_{2})$ are also the same with (4.27). Hence, we have proved the Proposition.
  2. (2)
    When $|I_{v_{3}}\cap J_{v_{3}}|=r_{v_{3}}-1,I_{v_{4}}=J_{v_{4}}$, the matrices representation for $\tau$ can be described as follows. Matrices $A_{u}$ for $u\neq v$ are reduced row/columns echelon forms with non-pivots entries vanishing and pivots columns/rows numbers are $I_{u},u\neq v$. The matrix $A_{v}$ is in the formula (4.22). Similar as the above situation, we still have
    $\mathbf{w}(\tau,\sigma_{1})=\lambda_{i_{1}}-\lambda_{j_{1}},\mathbf{w}(\tau,\sigma_{2})=\lambda_{j_{1}}-\lambda_{i_{1}}.$ (4.30)
    After the quiver mutation $\mu_{v}$, the matrices representations for $\varphi(\tau)$ can be described as follows. For $u\neq v,u\neq v_{3}$, matrices $A_{u}^{\prime}$ are exactly the same with $A_{u},\,u\neq v,\,u\neq v_{3}$. The matrices $A_{v_{3}}^{\prime}$ and $A_{v}^{\prime}$ are in the formula (4.17) and (4.18). The induced torus $(\mathbb{C}^{*})^{r_{v_{1}}}$ acts on $\tau$ as
    $t\cdot(A_{v_{3}}^{\prime},A_{v}^{\prime})=(A_{v_{3}}^{\prime}t^{-1},tA_{v}^{\prime}).$ (4.31)
@@
    $\mathbf{w}(\varphi(\tau),\varphi(\sigma_{1}))=\lambda_{i_{1}}-\lambda_{j_{1}},\,\mathbf{w}(\varphi(\tau),\varphi(\sigma_{2}))=\lambda_{j_{1}}-\lambda_{i_{1}}.$ (4.32)
    Comparing equations (4.30) and (4.32), we can find that the Proposition is true in this situation.
  3. (3)
@@ 4.6 Contributions of decorated graphs under quiver mutations
For each decorated graph $\vec{\Gamma}$, whose underling graph is $\Gamma$, the ingredients in Theorem 4.5 are weights of flags $\mathbf{w}(\tau,\sigma)$, the valency $val(v)$ of each vertex, and $h(\tau,d_{\tau})$ for each edge. Flags $(\tau,\sigma)\in\Upsilon$ are mapped to flags in $\varphi(\Upsilon)$, so $val(v)=val(\varphi(v))$. The weights $\mathbf{w}(\tau,\sigma)$ are preserved by Proposition 4.16.
Let $\tau$ be an edge, and $\sigma_{1},\sigma_{2}$ be two vertices contained in $\tau$. Let $E(\sigma_{1})=\{\tau_{1},\ldots,\tau_{h}\}$ be the set of edges containing the point $\sigma_{1}$. Without loss of generality, we assume that $\tau_{h}=\tau$. For each $i$, we have defined $a_{i}$ in Equation (4.4). We don’t have to compute all the $a_{i}$, but instead we only have to know the behavior of $a_{i}$ under quiver mutations. By Proposition 4.15, $E(\varphi(\sigma_{1}))=\{\varphi(\tau_{1}),\ldots,\varphi(\tau_{h})\}$, and $\varphi(\tau)$ contains points $\varphi(\sigma_{1})$ and $\varphi(\sigma_{2})$. By Proposition 4.16, weights are also preserved, so the Equation (4.4) still holds. We have
$\mathbf{w}(\varphi(\tau_{i}^{\prime}),\varphi(\sigma_{2}))=\mathbf{w}(\varphi(\tau_{i}),\varphi(\sigma_{1}))-a_{i}\mathbf{w}(\varphi(\tau),\varphi(\sigma_{1})).$ (4.33)
@@
Furthermore, the map $\varphi:\mathcal{Z}^{1}\rightarrow(\mathcal{Z}^{\prime})^{1}$ induces a map $\varphi_{*}:H_{2}(\mathcal{Z})\rightarrow H_{2}(\mathcal{Z}^{\prime})$. We claim that $\varphi_{*}$ coincides with $\phi_{*}$ in Proposition 4.10. In fact, suppose $\ell$ is a curve in $\mathcal{Z}$ corresponding to an edge of a graph in $\mathcal{Z}^{1}$. For an arbitrary $\gamma\in H^{2}(\mathcal{Z})$, we apply Atiyah-Bott localization formula [AB84] on the following two integrals $\int_{[\ell]}\gamma$ and $\int_{\varphi_{*}([\ell])}\phi(\gamma)$. Due to (4.9) and Proposition 4.16, we know the above two integrals are equal after comparing the graph contribution on both sides. Therefore, for an arbitrary $\gamma\in H^{2}(\mathcal{Z})$, we have
@@
By Proposition 4.10, $\varphi_{*}$ satisfies the same property (3.9). Due to the Poincare duality, we have $\varphi_{*}=\phi_{*}$. The proof of Theorem 4.1 is completed.
@@ 5 Proof of the cluster algebra conjecture for $A$-type
Until now, we have introduced the story of algebraic side (cluster algebra) and the geometric side (quantum cohomology). In this section we are going to establish a connection between them.
@@
$0\rightarrow S_{l}\rightarrow S_{k}\rightarrow S_{k}/S_{l}\rightarrow 0,$ (5.1)
@@
Since all relations in cluster algebra are from exchange relations (2.1), we need to prove that the map $\psi$ preserves cluster exchange relations.
@@
  1. (1)
    In the quantum cohomology ring of a flag variety $QH^{*}_{S}(Fl)[t]$, for $m>k\geq p\geq l\geq 0$, we have the following quantum cohomological cluster exchange relations.
@@
$x_{v}^{\prime}\tilde{x}_{v}=x_{v_{1}}^{\prime}x_{v_{4}}^{\prime}+x_{v_{2}}^{\prime}x_{v_{4}}^{\prime}.$ (5.4)
In order to prove that $\psi$ is a well-defined ring homomorphism, by the definition of $\psi$, we only have to prove that
@@
$+(-1)^{N_{k}-N_{k+1}+N_{p}-N_{p+1}}\xi_{p}\xi_{p+1}\xi_{k}^{-1}\xi_{k+1}^{-1}c_{t}^{S}(S_{m}/S_{k+1})*c_{t}^{S}(S_{p}/S_{l})\,,$ (5.6)
where the second term of the right hand side is exactly $\prod_{a=p+1}^{k}(-1)^{N_{a}+N_{a-1}}q_{a}c_{t}^{S}(S_{m}/S_{k+1})*c_{t}^{S}(S_{p}/S_{l})$ by letting $q_{a}=(-1)^{N_{a}-N_{a+1}}\xi_{a-1}\xi_{a+1}^{-1}$ Thus $\psi$ is a ring homomorphism.
The map $\psi$ is injective. Suppose $\alpha=f(x_{1},x_{2},\ldots,x_{n+1},\mathbf{x}^{\prime})\in\mathscr{A}_{n}$ satisfies $\psi(\alpha)=0$, we will check $\alpha=0$ in $\mathscr{A}_{n}$. By the Laurent phenomenon of cluster variables [FZ02, Theorem 3.1], any cluster variable can be expressed as a Laurent polynomial of in terms initial cluster variables. So there exist polynomial $P$ and monomial $Q$ in $\mathbb{Z}[x_{1},x_{2},\cdots,x_{n+1}]$ such that $Q(x_{1},\ldots,x_{n+1})\cdot\alpha=P(x_{1},\ldots,x_{n+1})$. Notice that $Q(x_{1},\ldots,x_{n+1})\neq 0$ and there is no zero divisor in $\mathscr{A}_{n}$. We can assume $\alpha=P(x_{1},\ldots,x_{n+1})$ without loss of generality. Suppose $\alpha$ is expressed as the following finite sum
@@
with $a_{\mathbf{i}}\in\mathbb{Q}$. We claim that $\psi(\alpha)=0$ implies $a_{\mathbf{i}}=0$ for arbitrary $\mathbf{i}$. To prove this, we introduce the following partial order $\geq$ in the set $\mathbb{Z}^{n+1}$: for $\mathbf{i}$ and $\mathbf{i}^{\prime}$ in $\mathbb{Z}^{n+1}$
@@
  4. (4)
    In the second author’s talk at Shanghai Jiao Tong University about this project, Professor Fomin informed us that they included a similar equation with Equation (1) in their book [FWZ21d], within the context of a different ring representation of the quantum cohomology ring of flag varieties.
In the remaining part of this section, we will prove that the Theorem 5.2 holds in $QH^{*}_{S}(Fl)[t]$. In Section 5.1-5.3, we will use the abelian/nonabelian correspondence for cohomology rings to study the cohomology and quantum cohomology rings of flag varieties. We want to emphasize that the ring representation is different from that in [CF95, CF99]. In Sections 5.4-5.5, we will give a detailed proof of the Theorem 5.2.

### v2

@@ 4.4 Behavior of one-dimensional torus-fixed orbits under quiver mutations
(25) $\varphi:\mathcal{Z}^{1}\rightarrow(\mathcal{Z}^{\prime})^{1}.$
@@
4.5 Behavior of weights under quiver mutations
@@
The main goal of this subsection is to investigate the behavior of weights of flags defined in (12) under quiver mutations.
@@ 4.5.1 Grassmannian case
Let $\sigma_{1}=I=\{i_{1}<\cdots<i_{r}\},\sigma_{2}=J=\{j_{1}<\cdots<j_{r}\}$ be two $S$-fixed points in $Gr(r,N)$ which, without loss of generality, satisfy $i_{1}\neq j_{1},i_{k}=j_{k}$ for $k\neq 1$. Let $\tau$ be the 1-dimensional $S$-orbit connecting the two points whose matrix representation is
(26) $\tau=\begin{bmatrix}\mathbf{0}&a&\mathbf{0}&b&\mathbf{0}\\ \mathbf{0}&\mathbf{0}&\mathbf{0}&\mathbf{0}&B\end{bmatrix}$
@@
(28) $\mathbf{w}(\tau,\sigma_{1})=c_{1}^{S}(T_{\sigma_{1}}\tau)=\lambda_{i_{1}}-\lambda_{j_{1}},\,\mathbf{w}(\tau,\sigma_{2})=c_{1}^{S}(T_{\sigma_{2}}\tau)=\lambda_{j_{1}}-\lambda_{i_{1}}.$
@@
(29) $\begin{bmatrix}\mathbf{0}&\mathbf{0}\\ b&\mathbf{0}\\ \mathbf{0}&\mathbf{0}\\ a&\mathbf{0}\\ \mathbf{0}&C\end{bmatrix}$
@@
(31) $\mathbf{w}(\varphi(\tau_{1}),\varphi(\sigma))=\lambda_{i_{1}}-\lambda_{j_{1}},\,\mathbf{w}(\varphi(\tau),\varphi(\sigma_{2}))=\lambda_{j_{1}}-\lambda_{i_{1}}.$
Compare equations (28) and (31), and one can find
$\mathbf{w}(\tau,\sigma_{1})=\mathbf{w}(\varphi(\tau_{1}),\varphi(\sigma)),\,\,\mathbf{w}(\tau,\sigma_{2})=\mathbf{w}(\varphi(\tau),\varphi(\sigma_{2}))\,.$
@@ 4.5.2 General case
$\mathbf{w}(\tau,\sigma)=\mathbf{w}(\varphi(\tau),\varphi(\sigma)).$
@@
  1. 1.
    When $I_{v_{3}}=J_{v_{3}}$, $I_{v_{4}}=J_{v_{4}}$, one matrix representative $(A_{u})_{u\in Q_{0}\backslash\{v_{f}\}}$ can be written as $A_{u},u\neq v$ are in reduced row/column echelon forms whose columns/rows numbers of pivots are $I_{u},u\neq v$. Meanwhile, the matrix $A_{v}$ can be written as the same formula as (26). We get the vertex $\sigma_{1}$ if $a=1,b=0$ and $\sigma_{2}$ is $a=0,b=1$. There is an induced torus action $(\mathbb{C}^{*})^{r_{v_{1}}}$ on the matrix $A_{v}$ as (27). Hence, we have $\mathbf{w}(\tau,\sigma_{1})=\lambda_{i_{1}}-\lambda_{j_{1}}$, and $\mathbf{w}(\tau,\sigma_{2})=\lambda_{j_{1}}-\lambda_{i_{1}}$.
    After the quiver mutation $\mu_{v}$, we then have $I_{v}^{\prime}\backslash J_{v}^{\prime}=\{j_{1}\}$ and $J_{v}^{\prime}\backslash I_{v}^{\prime}=\{i_{1}\}$. The matrices representatives $A_{u}^{\prime},\,u\neq v$ are the same with $A_{u},\,u\neq v$, but $A_{v}^{\prime}$ is changed to the formula (29), with the induced torus action $(\mathbb{C}^{*})^{r_{v_{1}}}$ the same with (30). One can easily find that the weights of the flag at the vertex $\varphi(\sigma_{1})$ and $\varphi(\sigma_{2})$ are also the same with (31). Hence, we have proved the Proposition.
  2. 2.
    When $|I_{v_{3}}\cap J_{v_{3}}|=r_{v_{3}}-1,I_{v_{4}}=J_{v_{4}}$, the matrices representation for $\tau$ can be described as follows. Matrices $A_{u}$ for $u\neq v$ are reduced row/columns echelon forms with non-pivots entries vanishing and pivots columns/rows numbers are $I_{u},u\neq v$. The matrix $A_{v}$ is in the formula (26). Similar as the above situation, we still have
    (32) $\mathbf{w}(\tau,\sigma_{1})=\lambda_{i_{1}}-\lambda_{j_{1}},\mathbf{w}(\tau,\sigma_{2})=\lambda_{j_{1}}-\lambda_{i_{1}}.$
    After the quiver mutation $\mu_{v}$, the matrices representations for $\varphi(\tau)$ can be described as follows. For $u\neq v,u\neq v_{3}$, matrices $A_{u}^{\prime}$ are exactly the same with $A_{u},\,u\neq v,\,u\neq v_{3}$. The matrices $A_{v_{3}}^{\prime}$ and $A_{v}^{\prime}$ are in the formula (23) and (24). The induced torus $(\mathbb{C}^{*})^{r_{v_{1}}}$ acts on $\tau$ as
    $t\cdot(A_{v_{3}}^{\prime},A_{v}^{\prime})=(A_{v_{3}}^{\prime}t^{-1},tA_{v}^{\prime}).$
@@
    (33) $\mathbf{w}(\varphi(\tau),\varphi(\sigma_{1}))=\lambda_{i_{1}}-\lambda_{j_{1}},\,\mathbf{w}(\varphi(\tau),\varphi(\sigma_{2}))=\lambda_{j_{1}}-\lambda_{i_{1}}.$
    Comparing equations (32) and (33), we can find that the Proposition is true in this situation.
  3. 3.
@@ 4.6 Contributions of decorated graphs under quiver mutations
For each decorated graph $\vec{\Gamma}$, whose underlying graph is $\Gamma$, the ingredients in Theorem 4.5 are weights of flags $\mathbf{w}(\tau,\sigma)$, the valency $val(v)$ of each vertex, and $h(\tau,d_{\tau})$ for each edge. Flags $(\tau,\sigma)\in\Upsilon$ are mapped to flags in $\varphi(\Upsilon)$, so $val(v)=val(\varphi(v))$. The weights $\mathbf{w}(\tau,\sigma)$ are preserved by Proposition 4.16.
Let $\tau$ be an edge, and $\sigma_{1},\sigma_{2}$ be two vertices contained in $\tau$. Let $E(\sigma_{1})=\{\tau_{1},\ldots,\tau_{h}\}$ be the set of edges containing the point $\sigma_{1}$. Without loss of generality, we assume that $\tau_{h}=\tau$. For each $i$, we have defined $a_{i}$ in Equation (13). We don’t have to compute all the $a_{i}$, but instead we only have to know the behavior of $a_{i}$ under quiver mutations. By Proposition 4.15, $E(\varphi(\sigma_{1}))=\{\varphi(\tau_{1}),\ldots,\varphi(\tau_{h})\}$, and $\varphi(\tau)$ contains points $\varphi(\sigma_{1})$ and $\varphi(\sigma_{2})$. By Proposition 4.16, weights are also preserved, so the Equation (13) still holds. We have
$\mathbf{w}(\varphi(\tau_{i}^{\prime}),\varphi(\sigma_{2}))=\mathbf{w}(\varphi(\tau_{i}),\varphi(\sigma_{1}))-a_{i}\mathbf{w}(\varphi(\tau),\varphi(\sigma_{1})).$
@@
Furthermore, the map $\varphi:\mathcal{Z}^{1}\rightarrow(\mathcal{Z}^{\prime})^{1}$ induces a map $\varphi_{*}:H_{2}(\mathcal{Z})\rightarrow H_{2}(\mathcal{Z}^{\prime})$. We claim that $\varphi_{*}$ coincides with $\phi_{*}$ in Proposition 4.10. In fact, suppose $\ell$ is a curve in $\mathcal{Z}$ corresponding to an edge of a graph in $\mathcal{Z}^{1}$. For an arbitrary $\gamma\in H^{2}(\mathcal{Z})$, we apply the Atiyah-Bott localization formula [ATIYAH19841] on the following two integrals $\int_{[\ell]}\gamma$ and $\int_{\varphi_{*}([\ell])}\phi(\gamma)$. Due to (17) and Proposition 4.16, we know the above two integrals are equal after comparing the graph contribution on both sides. Therefore, for an arbitrary $\gamma\in H^{2}(\mathcal{Z})$, we have
@@
By Proposition 4.10, $\varphi_{*}$ satisfies the same property (8). Due to the Poincaré duality, we have $\varphi_{*}=\phi_{*}$. The proof of Theorem 4.1 is completed.
@@ 5 Proof of the cluster algebra conjecture for A-type
Until now, we have introduced the story of the algebraic side (cluster algebra) and the geometric side (quantum cohomology). In this section, we are going to establish a connection between them.
@@
(34) $0\rightarrow S_{l}\rightarrow S_{k}\rightarrow S_{k}/S_{l}\rightarrow 0,$
@@
Since all relations in cluster algebra are from exchange relations (2), we need to prove that the map $\psi$ preserves cluster exchange relations.
@@
  1. 1.
@@
$x_{v}^{\prime}\tilde{x}_{v}=x_{v_{1}}^{\prime}x_{v_{4}}^{\prime}+x_{v_{2}}^{\prime}x_{v_{4}}^{\prime}.$
In order to prove that $\psi$ is a well-defined ring homomorphism, we only have to prove that
@@
$+(-1)^{N_{k}-N_{k+1}+N_{p}-N_{p+1}}\xi_{p}\xi_{p+1}\xi_{k}^{-1}\xi_{k+1}^{-1}c_{t}^{S}(S_{m}/S_{k+1})*c_{t}^{S}(S_{p}/S_{l})\,,$
By letting $q_{a}=(-1)^{N_{a}-N_{a+1}}\xi_{a-1}\xi_{a+1}^{-1}$, the second term of the right-hand side is exactly $\prod_{a=p+1}^{k}(-1)^{N_{a}+N_{a-1}}q_{a}c_{t}^{S}(S_{m}/S_{k+1})*c_{t}^{S}(S_{p}/S_{l})$. Thus $\psi$ is a ring homomorphism.
The map $\psi$ is injective. Suppose $\alpha=f(x_{1},x_{2},\ldots,x_{n+1},\mathbf{x}^{\prime})\in\mathscr{A}_{n}$ is in the kernel of $\psi$. We will prove that $\alpha=0$ in $\mathscr{A}_{n}$. By the Laurent phenomenon of cluster variables [clusteralg:1, Theorem 3.1], any cluster variable can be expressed as a Laurent polynomial of initial cluster variables. Therefore, there exists a polynomial $P$ and a monomial $Q$ in $\mathbb{Z}(x_{1},x_{2},\cdots,x_{n+1})$ such that $Q(x_{1},\ldots,x_{n+1})\cdot\alpha=P(x_{1},\ldots,x_{n+1})$. Since there is no zero divisor in $\mathscr{A}_{n}$, we only have to prove that $P=0$ in $\mathscr{A}_{n}$. We can assume $\alpha=P(x_{1},\ldots,x_{n+1})$ without loss of generality. Suppose $\alpha$ is expressed as the following finite sum
@@
with $a_{\mathbf{i}}\in\mathbb{Q}$. We claim that $\psi(\alpha)=0$ implies $a_{\mathbf{i}}=0$ for arbitrary $\mathbf{i}$. We introduce the following partial order $\geq$ in the set $\mathbb{Z}^{n+1}$: for $\mathbf{i}$ and $\mathbf{i}^{\prime}$ in $\mathbb{Z}^{n+1}$
@@
  4. 4.
    In the second author’s talk at Shanghai Jiao Tong University about this project, Professor Fomin informed us that they included a similar equation with Equation (1) in their book [fomin:cluster4-5], within the context of a different ring representation of the quantum cohomology ring of flag varieties.
In the remaining part of this section, we will prove that the Theorem 5.2 holds in $QH^{*}_{S}(Fl)[t]$. In Section 5.1-5.3, we will use the abelian/nonabelian correspondence for cohomology rings to study the cohomology and quantum cohomology rings of flag varieties. We want to emphasize that the ring representation is different from that in [flag:quantumcohomology, qcohflag:CF99]. In Sections 5.4-5.5, we will give a detailed proof of the Theorem 5.2.